4x-y= -1
2x+4y=13

Answer by skerry(1)

4x-y = -1 ...................... equation(1)
2x+4y = 13 .......................equation(2)
to apply the elimination method you multiply equation 2 by 2.
4x+8y = 26 .......................equation(3)
adding equation 1 and 3 in order to eliminate x you have:
4x-y = -1....................(1)
4x+8y = 26....................(3)
equation. (1) + (3)
-9y = -27
divide both sides by -9
therefore y = 3
replace y as 3 in equation. (1)
4x- 3= -1
4x = 3-1
4x = 2
divide both sides by 4 to x = 1/2
therefore y = 3 and x = 1/2 (Answer)